Solve using the multiplication principle first. Then add. Then show your check for both equations.
-x-y=8 equation (1)
2x-y=-1 equation (2)
I think I have the correct answers. x=-3 and y=-5. My problem is my check on the first equation -x-y=8. I cannot get it to check. The second equation checks. 2x-y=-1: 2*3-(-5)=-1. Can you show me how to check -x-y=8 please. I have spent an hour trying to get this to check and it won't. This algebra area we are doing is System of equations the elimination method and we have to do the multiplication principle first, then add, then do our checks for each equation. I have to turn this in first thing in the morning and I am frustrated at this point because I cannot make the first problem check. Please help.

Found 2 solutions by solver91311, ewatrrr:
Answer by solver91311(24713)



You say the solution set is the ordered pair (-3,-5), so:







Looks like it checks to me.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by ewatrrr(24785)


Hi
Solve using the multiplication principle first. Then add.
-x-y=8 equation (1)
2x-y=-1 equation (2)
-2x-2y=16 |multiplying 1st Eq thru by 2 and then adding to eliminate x
2x-y =-1
-3y = 15
y = -5 and
2x = y-1 = -6;
x = -3
CHECKING our Answer***
3 + 5 = 8 |this checks... -(-3) -(-5) = 3 + 5 = 8
-6+5 = -1